This is a problem from a quiz of mine and I have the answer but I don't know how to get it. Jimmy can only haverage 40 mph when he fights Friday afternoon traffic to drive the 120 miles to his vacation home. Coming home on Sunday morning, however, he is able to average 60 mph. What is his average speed for the round trip? Thanks!
Answer by venugopalramana(3286)

This is a problem from a quiz of mine and I have the answer but I don't know how to get it. Jimmy can only haverage 40 mph when he fights Friday afternoon traffic to drive the 120 miles to his vacation home.
D=120 MILES
S1=40 MPH
T1=D/S=120/40=3 HRS.
Coming home on Sunday morning, however, he is able to average 60 mph.
S2=60 MPH
T2=120/60=2 HRS.
HENCE FOR ROUND TRIP,WE HAVE
TOTAL DISTANCE TRAVELLED=TD=120+120
TOTAL TIME TAKEN =TT=T1+T2=3+2=5 HRS.
HENCE AVERAGE SPEED =TD/TT=240/5=48 MPH
What is his average speed for the round trip? Thanks!